OpenStudy (anonymous):

2x – 2 < –12 or 2x + 3 > 7

OpenStudy (anonymous):

idk what the quiston is tho, are you trying to single out x from both sides?

OpenStudy (anonymous):

if so i can help @leedel_15

OpenStudy (anonymous):

2x-2<12 +2 +2 2x<10 /2 /2 x<5

OpenStudy (anonymous):

2x+3>7 -3 -3 2x>7 /2 /2 x>3,5

OpenStudy (anonymous):

\(\large 2x-2<12 \Rightarrow 2x<12+2=14 \Rightarrow x<\frac{14}{2}=7\\\large 2x+3>7\Rightarrow 2x>7-3=4\Rightarrow x>2\)